Gary has a bag containing 15 trucks, 3 cars, 8 trains, and 6 airplanes. If he randomly chooses one piece from the bag, what is the probability that he will grab a truck or a car?
Would it be 15/32 for a truck and 3/32 for a car

Answer by stanbon(75887)

Gary has a bag containing 15 trucks, 3 cars, 8 trains, and 6 airplanes. If he randomly chooses one piece from the bag, what is the probability that he will grab a truck or a car?
Would it be 15/32 for a truck and 3/32 for a car
-----
Yes, but this is an "OR" problem so you have to add the two
probabilities to get 15/32 + 3/32 = 18/32 = 9/16
===========================
Cheers,
Stan H.